You turned the knob of a 194.22 cm by 103.86 cm door (with a mass of 9.05 kg) and pushed it with a perpendicular force of 5.18N to open it. What is the approximate torque on the door in (N⋅m)?

1
Expert's answer
2021-11-29T11:44:30-0500

"\tau=F\cdot d=5.18\cdot 1.0386=5.38\ (N\cdot m)" . Answer